OpenStudy (ambermarie151):

Factor the polynomial. -3x^4-9x^3+15x^2

OpenStudy (campbell_st):

well start by identifying the highest common factor in each term... any thoughts..?

OpenStudy (ambermarie151):

So like x4 +x3+x2

OpenStudy (campbell_st):

close, each term has a coefficient that is a multiple of 3 each term has a veriable in x so the highest common factor is 3x^2 or even -3x^2 so to start the factorisation it can be written as \[-3x^2(~~~~~~~~~~~~~~~)\] there will be 2 terms inside the brackets, what will they be..?

OpenStudy (campbell_st):

oops... that should read, there will be 3 terms inside the brackets... sorry about the typo

OpenStudy (ambermarie151):

would be like x -9 +15

OpenStudy (campbell_st):

not quite... what can you simplify these fractions. \[\frac{-3x^4}{-3x^2}\] and \[\frac{-9x^3}{-3x^2}\] then \[\frac{15x^2}{-3x^2}\]

OpenStudy (ambermarie151):

So it will be 1+3-5